Sol Summary – March 6th

Crew 263 Sol Summary 06-03-2022

Sol: 13

Summary Title: Gypsum and Suspicious Gas Smell

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: Yesterday, we decided to start the sport session 30 minutes later than usual to have 30 minutes of bonus sleep. We will probably keep this timing for the rest of the mission as it made us feel more energetic. We were happy to see that there was no snow this morning, allowing us to do our EVA as planned, in the North, between Tank Wash and White Moons. We were not expecting gypsum as we saw it, but took samples to analyse in the Hab. After talking with Shannon about it and a quick LIBS analysis, we found out they were indeed Gypsum, which was quite satisfying. It was our first time in this area and we found it beautiful.

In the afternoon, there wasn’t a lot of scientific activities. We thus had to cheat boredom: some of us did colour painting supervised by our specialist, Marine, Nicolas and Mathéo did a video tour of the station, and I knitted. We had a little scare in the evening: a very strong smell of gas spread in the Hab, coming from the vents. Shannon radio-assisted us in the search for a leak: none was detected and 15 minutes later we could not smell it anymore.

Look Ahead Plan: We were very excited to plan Tuesday’s EVA that will probably be the most relevant one for the research team back in Toulouse that trained Marine and Lea to Geology and LIBS analysis. As it will be a very long and surely exhausting EVA, we planned tomorrow’s EVA accordingly, so that Tuesday’s EVA team rests the day before. Tomorrow I’ll turn 25: I am sure the Crew will make it a special day!

Anomalies in work: Gas odour in the Hab in the evening. No leak detected, and odour stopped.

Weather: Sunny in the morning, cloudy and very windy in the afternoon.

Crew Physical Status: Good.

EVA: EVA 17 in the morning. No issue to report.

Reports to be filed: Sol Summary, EVA Report (n°17), EVA Request (n°18), Operations Report, Journalist Report, HSO Report, Scientist Report.

Support Requested: No support requested.